OpenStudy (anonymous):

WILL MEDAL. what is the radius of the circle given by the following equation. x^2/10+(y-2)^2/10=1 a) 30 b) sqrt10 c) 5 d) 1

OpenStudy (anonymous):

\[\frac{x^2}{10}+\frac{(y-2)^2}{10}=1\\ x^2+(y-2)^2=10\] now you can read off the answer

OpenStudy (anonymous):

general form is \[(x-h)^2+(y-k)^2=r^2\] for a circle with center \((h,k)\) and radius \(r\) you have \[x^2+(y-2)^2=10\]

OpenStudy (anonymous):

so the answer would be b? @satellite73

OpenStudy (anonymous):

yes